# Comparison of modularity-based approaches for nodes clustering in binary hypergraphs

### 25 Jan 2024 v1

@@ Abstract
We conducted a comparative analysis of the performance of modularity-based methods for clustering nodes in binary hypergraphs. Statistical analysis and node clustering in hypergraphs constitute an emerging topic suffering from a lack of standardization. In contrast to the case of graphs, the concept of nodes’ community in hypergraphs is not unique and encompasses various distinct situations.
To address this, we begin by presenting, within a unified framework, the various hypergraph modularity criteria proposed in the literature, emphasizing their differences and respective focuses. Subsequently, we provide an overview of the state-of-the-art codes available to maximize hypergraph modularities for detecting node communities in binary hypergraphs. Through exploration of various simulation settings with controlled ground truth clustering, we offer a comparison of these methods using different quality measures, including true clustering recovery, running time, (local) maximization of the objective, and the number of clusters detected.
@@ 1 Introduction
With the emergence of hypergraph datasets to model higher-order interactions, the question of nodes clustering and, more specifically, the detection of communities in hypergraphs arises. In the context of graphs, the seminal paper by Newman and Girvan (2004) introduced the concept of modularity (commonly known as the Newman-Girvan modularity), paving the way for a flourishing literature on community detection in networks. In the context of hypergraphs, the past few years have witnessed the surge of modularity-based proposals for hypergraph community detection. One of the first challenges is to propose a modularity criterion that measures the extent to which a hypergraph is composed of communities. This raises a more fundamental question: What is a community of nodes in a hypergraph? While in the context of graphs, a community is simply a set of nodes with more within-cluster interactions than between-clusters ones, generalizing that concept to hypergraphs is not immediate. As hypergraph interactions have a heterogeneous size (i.e., the number of nodes they contain), a primary issue is whether one should weigh the links with respect to (wrt) their sizes and put more emphasis on larger hyperedges (see Figure 1 for an illustration). Consequently, various modularity criteria have recently emerged in the literature.
@@
In this article, we explore the current state-of-the-art and challenges posed by modularity-based community detection methods in binary hypergraphs. In the context of graphs, Yang et al. (2016) propose a comparative analysis of community detection algorithms for undirected and binary graphs. In the same vein, we here restrict our attention to modularity-based methods whose performances for community detection in binary hypergraphs are compared. The methodology is described in Section 2. After introducing general notation, we first present a reformulated version of the different hypergraph modularities existing in the literature (Section 2.2). The goal of this reformulation is to facilitate the comparison of concepts introduced independently from each other and never fully connected before. To be a valuable concept, a hypergraph modularity should come with a (local and/or heuristic) maximization algorithm that outputs a node clustering. Available implementations of such algorithms are presented in Section 2.3. To compare the different modularities and maximization algorithms, it is mandatory to work with synthetic datasets where ground truth clustering is known and hypergraph statistics can be controlled. While in the graph context, recent years have seen the emergence of benchmark datasets for such a task, as for instance the Lancichinetti-Fortunato-Radicchi benchmark graph (LFR, Lancichinetti et al., 2008) used in Yang et al. (2016), there is yet no such benchmark for hypergraphs. We thus rely on several models for generating synthetic modular hypergraphs, described in Section 2.4. Then Section 3 describes our experiments: which scenarios have been explored in each model generating method (Section 3.1) and quality assessment through the lens of different measures, namely true clustering recovery, running time, (local) maximization of the objective, number of clusters detected (Section 3.2). All the results are presented in Section 4 and a discussion follows in Section 5. The scripts to reproduce the experiments are available online (see details and links in Section 6).
@@ 2.1 General notation and definitions
From a (weighted) hypergraph $H=(V,\mathcal{E})$ we may construct its clique reduction graph $G^{\textrm{clique}}=(V,E)$. This graph has the same set of nodes $V$ than the hypergraph and every hyperedge $e\in\mathcal{E}$ in the hypergraph is reduced into a complete clique in the graph. In other words, for any hyperedge $e\in\mathcal{E}$ with size $|e|\geq 2$ and for any pair of incident nodes $u,v\in e$, the graph $G^{\textrm{clique}}$ contains the edge $\{u,v\}\in E$ and only edges obtained in this way are contained in $E$. The (weighted) adjacency matrix $A^{\textrm{clique}}$ of the clique reduction graph satisfies $A^{\textrm{clique}}=H\mathrm{diag}(w)H^{\intercal}$, where $H$ is the incidence matrix of the hypergraph and $\mathrm{diag}(w)$ is the diagonal matrix induced by the vector of hyperedge weights $w$. In general, self-loops are removed from $A^{\textrm{clique}}$ and $A^{\textrm{clique}}_{uu}$ is set to 00 for any $u\in V$. This can be done directly by setting $A^{\textrm{clique}}=H\mathrm{diag}(w)H^{\intercal}-D_{V}$ where $D_{V}$ is the diagonal matrix of node degrees $(\deg_{H}(v))_{v\in V}$.
@@ 2.2 Modularities in hypergraphs
For instance, a $s$-tuple of nodes with $s=7$ that are clustered by a partition $\mathcal{C}$ into the parts $\{v_{1}\};\{v_{2},v_{4}\};\{v_{3}\};\{v_{5},v_{6},v_{7}\}$ induces the partition vector $\mathbf{p}=(3,2,1,1)$. The symmetric modularity from Chodrow et al. (2021) will thus account for the different clusters counts that compose a hyperedge, treating all the clusters in an exchangeable way. We present the details of this modularity in Section A from the Appendix. Then, the authors consider particular cases of their general symmetric modularity, relying on specific forms of the affinity function $\Omega$ (see Table 1 in that reference). However, an implementation of the algorithm for optimising the induced specific modularities is available only for the all-or-nothing affinity function on which we focus now.
@@
Two extreme cases are represented by the modularities $Q^{\textrm{w-clique}}$ and $Q^{\textrm{strict}}$, the former being less stringent than the latter. Whenever a hyperedge is split by the partition $\mathcal{C}$ into different clusters, it will be ignored by $Q^{\textrm{strict}}$ but as soon as this hyperedge contains at least 2 nodes in the same cluster, the modularity $Q^{\textrm{w-clique}}$ will account for it. The weakness in $Q^{\textrm{w-clique}}$ lies in that the exact composition of each hyperedge in nodes falling into the different clusters is captured only through pairs of nodes. The modularity $Q^{\textrm{wsc}}$ represents a compromise between the 2 previous extremes: it accounts for homogeneous hyperedges, namely hyperedges such that (at least) half of their nodes fall into cluster that becomes a majority cluster. In particular, Kamiński et al. (2021) argue that the hyper-parameters $w_{s,c}$ may be chosen so that $Q^{\textrm{wsc}}$ well approximates $Q^{\textrm{w-clique}}$ because contributions in the latter from parts that contain at most $s/2$ vertices may often be neglected. Finally, the modularity $Q^{\textrm{aon}}$ is as strict as $Q^{\textrm{strict}}$ and focuses on hyperedges with nodes split into a unique cluster by the partition $\mathcal{C}$. As already stressed, the major difference between $Q^{\textrm{strict}}$ and $Q^{\textrm{aon}}$ lies in that the latter, while summing similar differences as the former, weights differently each terms in those differences (with weights adaptive to the data, as they are estimated from these).
@@ 2.3 Modularity maximization methods
Kumar et al. (2020) propose to maximize their modularity $Q^{\textrm{w-clique}}$ relying on the popular and fast Louvain algorithm for graphs (Blondel et al., 2008). More precisely, they do not simply apply Louvain algorithm on the graph $G^{\textrm{w-clique}}$ but rather propose an Iteratively Reweighted Modularity Maximization (IRMM) algorithm where they iteratively apply Louvain on a weighted clique reduction graph, and compute new hyperedge weights (see Algorithm 1 in Kumar et al., 2020). The hyperedge re-weighting step puts a larger weight on hyperedges which are cut into more unbalanced partition vectors by the current partition $\mathcal{C}$. For example, a size-$s$ hyperedge cut into the partition vector $\mathbf{p}=(s-1,1)$ (meaning a unique node falls in a cluster different from the majority one) is much more unbalanced than another one cut into the partition vector $\mathbf{p}=(s/2,s/2)$ (namely half of the nodes belong to a first cluster, the other half belonging to a second cluster) and thus gets a larger weight (see Figure 1 in Kumar et al., 2020). By getting a larger weight, it is more likely that the unique node in this hyperedge will join the majority cluster at Louvain’s next step. The function hmod.kumar() implements the IRMM algorithm.
@@
Kamiński et al. (2019a) propose a Clauset-Newman-Moore like (CNM-like) algorithm to maximize $Q^{\textrm{strict}}$ (see Clauset et al., 2004, for the original CNM algorithm). Starting with partition $\mathcal{C}^{\text{own}}$ where each node is in its own part, this algorithm iterates over the set of hyperedges that are split into more than 2 clusters by the current partition, trying to merge all the parts it touches and looking for a modularity improvement. More precisely, the algorithm has two versions. In the first one, a loop over all hyperedges is taken, so that at each step all hyperedges are searched and evaluated for merging. In the second one, a stochastic approach is taken which evaluates at each step just one randomly chosen hyperedge (see Algorithm 1 in Kamiński et al., 2019a, for more details).  The stochastic version is computationally less expensive, especially for larger hypergraphs; however it requires to set a maximal number of iterations. In what follows, we choose to set this number of iterations to twice the total number of hyperedges. The implementation is available from Kamiński et al. (2019b), in a mix of Python and Julia files. More precisely, a script strictModularity.py contains a “quick” Python implementation that should work on small datasets only, while a Julia function find_comms() is more generally provided to perform the CNM-like algorithm. We rely on the latter in our experiments.
@@ 2.4 Synthetic models for binary and modular hypergraphs
To compare the different modularity-based approaches for clustering hypergraphs nodes, it is mandatory to rely on simulations of modular hypergraphs where ground-truth clusters are known. As mentioned earlier, there is no single standard method for generating modular hypergraphs, and, to our knowledge, there are two main approaches. The first approach is based on hypergraph stochastic block models, with several variants proposed in the literature. The second approach involves a generalization of the LFR model for graphs (Lancichinetti et al., 2008). We chose to consider two variants of the first approach and the only one that we are aware of in the second approach. A summary of these models is given in Table 2. We highlight the similarities and differences between those different generating models and the characteristics of the hypergraphs generated by those approaches.
In all those models, we fix a number of nodes $n$, either fix or generate a true number of clusters $K$ (that might depend on $n$) as well as a true partition of the nodes $\mathcal{C}^{\text{true}}=(C_{1}^{\text{true}},\dots,C_{K}^{\text{true}})$ and a maximal size $S$ of hyperedges.
@@ 3.1 Scenarios
We start with “base case” scenarios, called scenarios A, that we declined in the 3 different generating models (HSBM, DCHSBM and h-ABCD). Then we explore other scenarios (called B to F and Z), simulated under the most convenient model to do so, and in which we modify only one characteristic at a time wrt the base case. Each scenario is composed of sub-cases with different samples sizes, comprising in general cases 1 to 6 corresponding to $n\in\{50,100,150,200,500,1000\}$. Moreover, in each scenario explored, we randomly generated 25 hypergraphs. Table 3 gives a summary of the scenarios considered and the empirical characteristics of the 25 hypergraphs generated for each of them. In this table, each simulation is summarized through its differing characteristic wrt the base case (namely, scenarios A). For example, ScenB-DCHSBM is a simulation of hypergraphs less sparse than the base case.
@@
We started from a first series of scenarios, called scenarios A, that play the role of a reasonable sparse case for the methods to work. To explore the robustness of our conclusions, these scenarios are presented under the 3 different generating models (HSBM, DCHSBM and h-ABCD) relying on similar settings for sample size $n$ and number of hyperedges $|\mathcal{E}|$. We set the numbers of hyperedges such that they grow linearly with the number of nodes $n$ (sparse setting). We generated $K=3$ clusters with equal size or probability (depending on the generating model) and the maximum hyperedge size $S=3$. The ratio $|\mathcal{E}_{2}|/|\mathcal{E}|=|\mathcal{E}_{2}|/(|\mathcal{E}_{2}|+|\mathcal{E}_{3}|)$ is set to 0.7 (on average) to reflect the fact that we expect larger sizes hyperedges to be less frequent than smaller-sizes ones. The within-cluster over between-cluster hyperedge ratio is constant wrt size $s\in\{2,3\}$ and set to $\rho_{s}=1.7$ (either exactly or on average), in order to obtain modular hypergraphs.

### 04 Mar 2024 v2

# Comparison of modularity-based approaches for nodes clustering in hypergraphs

@@ Abstract
Statistical analysis and node clustering in hypergraphs constitute an emerging topic suffering from a lack of standardization. In contrast to the case of graphs, the concept of nodes’ community in hypergraphs is not unique and encompasses various distinct situations. In this work, we conducted a comparative analysis of the performance of modularity-based methods for clustering nodes in binary hypergraphs.
To address this, we begin by presenting, within a unified framework, the various hypergraph modularity criteria proposed in the literature, emphasizing their differences and respective focuses. Subsequently, we provide an overview of the state-of-the-art codes available to maximize hypergraph modularities for detecting node communities in hypergraphs. Through exploration of various simulation settings with controlled ground truth clustering, we offer a comparison of these methods using different quality measures, including true clustering recovery, running time, (local) maximization of the objective, and the number of clusters detected.
@@ 1 Introduction
With the emergence of hypergraph datasets (see for e.g. Lee et al., 2021) to model higher-order interactions, the question of nodes clustering and, more specifically, the detection of communities in hypergraphs arises. In the context of graphs, the seminal paper by Newman and Girvan (2004) introduced the concept of modularity (commonly known as the Newman-Girvan modularity), paving the way for a flourishing literature on community detection in networks. In the context of hypergraphs, the past few years have witnessed the surge of modularity-based proposals for hypergraph community detection. One of the first challenges is to propose a modularity criterion that measures the extent to which a hypergraph is composed of communities. This raises a more fundamental question: What is a community of nodes in a hypergraph? While in the context of graphs, a community is simply a set of nodes with more within-cluster interactions than between-clusters ones, generalizing that concept to hypergraphs is not immediate. As hypergraph interactions have a heterogeneous size (i.e., the number of nodes they contain), a primary issue is whether one should weigh the links with respect to (wrt) their sizes and put more emphasis on larger hyperedges (see Figure 1 for an illustration). Consequently, various modularity criteria have recently emerged in the literature.
@@
In this article, we explore the current state-of-the-art and challenges posed by modularity-based community detection methods in binary hypergraphs. In the context of graphs, Yang et al. (2016) propose a comparative analysis of community detection algorithms for undirected and binary graphs. In the same vein, we here restrict our attention to modularity-based methods whose performances for community detection in binary hypergraphs are compared. The methodology is described in Section 2. After introducing general notation, we first present a reformulated version of the different hypergraph modularities existing in the literature (Section 2.2). The goal of this reformulation is to facilitate the comparison of concepts introduced independently from each other and never fully connected before. To be a valuable concept, a hypergraph modularity should come with a (local and/or heuristic) maximization algorithm that outputs a node clustering. Available implementations of such algorithms are presented in Section 2.3. To compare the different modularities and maximization algorithms, it is mandatory to work with synthetic datasets where ground truth clustering is known and hypergraph statistics can be controlled. While in the graph context, recent years have seen the emergence of benchmark datasets for such a task, as for instance the Lancichinetti-Fortunato-Radicchi benchmark graph (LFR, Lancichinetti et al., 2008) used in Yang et al. (2016), there is yet no such benchmark for hypergraphs. We thus rely on several models for generating synthetic modular hypergraphs, described in Section 2.4. Then Section 3 describes our experiments: which scenarios have been explored in each model generating method (Section 3.1) and quality assessment through the lens of different measures, namely true clustering recovery, running time, (local) maximization of the objective and the number of clusters detected (Section 3.2). All the results are presented in Section 4 and a discussion follows in Section 5. The scripts to reproduce the experiments are available online (see details and links in Section 6).
To conclude this introduction, we mention that there are other methods to cluster the nodes of a hypergraph, such as spectral clustering approaches (Ghoshdastidar and Dukkipati, 2017; Chodrow et al., 2023) or model-based methods (Brusa and Matias, 2022a; Ruggeri et al., 2023). It is also possible to cluster hyperedges instead of nodes (Ng and Murphy, 2022). However, our focus in this work is on clustering nodes through modularity-based methods.
@@ 2.1 General notation and definitions
From a (weighted) hypergraph $H=(V,\mathcal{E})$ we may construct its clique reduction graph $G^{\textrm{clique}}=(V,E)$. This graph has the same set of nodes $V$ as the hypergraph and every hyperedge $e\in\mathcal{E}$ in the hypergraph is reduced into a complete clique in the graph. In other words, for any hyperedge $e\in\mathcal{E}$ with size $|e|\geq 2$ and for any pair of incident nodes $u,v\in e$, the graph $G^{\textrm{clique}}$ contains the edge $\{u,v\}\in E$ and only edges obtained in this way are contained in $E$. The (weighted) adjacency matrix $A^{\textrm{clique}}$ of the clique reduction graph satisfies $A^{\textrm{clique}}=H\mathrm{diag}(w)H^{\intercal}$, where $H$ is the incidence matrix of the hypergraph and $\mathrm{diag}(w)$ is the diagonal matrix induced by the vector of hyperedge weights $w$. In general, self-loops are removed from $A^{\textrm{clique}}$ and $A^{\textrm{clique}}_{uu}$ is set to 00 for any $u\in V$. This can be done directly by setting $A^{\textrm{clique}}=H\mathrm{diag}(w)H^{\intercal}-D_{V}$ where $D_{V}$ is the diagonal matrix of node degrees $(\deg_{H}(v))_{v\in V}$.
@@ 2.2 Modularities in hypergraphs
For instance, a $s$-tuple of nodes with $s=7$ that are clustered by a partition $\mathcal{C}$ into the parts $\{v_{1}\};\{v_{2},v_{3}\};\{v_{4}\};\{v_{5},v_{6},v_{7}\}$ induces the partition vector $\mathbf{p}=(3,2,1,1)$. The symmetric modularity from Chodrow et al. (2021) will thus account for the different clusters counts that compose a hyperedge, treating all the clusters in an exchangeable way. We present the details of this modularity in Section A from the Appendix. Then, the authors consider particular cases of their general symmetric modularity, relying on specific forms of the affinity function $\Omega$ (see Table 1 in that reference). However, an implementation of the algorithm for optimising the induced specific modularities is available only for the all-or-nothing affinity function on which we focus now.
@@
Two extreme cases are represented by the modularities $Q^{\textrm{w-clique}}$ and $Q^{\textrm{strict}}$, the former being less stringent than the latter. Whenever a hyperedge is split by the partition $\mathcal{C}$ into different clusters, it will be ignored by $Q^{\textrm{strict}}$ but as soon as this hyperedge contains at least 2 nodes in the same cluster, the modularity $Q^{\textrm{w-clique}}$ will account for it. The weakness in $Q^{\textrm{w-clique}}$ lies in that the exact composition of each hyperedge in nodes falling into the different clusters is captured only through pairs of nodes. The modularity $Q^{\textrm{wsc}}$ represents a compromise between the 2 previous extremes: it accounts for homogeneous hyperedges, namely hyperedges such that (at least) half of their nodes fall into a cluster that becomes a majority cluster. In particular, Kamiński et al. (2021) argue that the hyper-parameters $w_{s,c}$ may be chosen so that $Q^{\textrm{wsc}}$ well approximates $Q^{\textrm{w-clique}}$ because contributions in the latter from parts that contain at most $s/2$ vertices may often be neglected. Finally, the modularity $Q^{\textrm{aon}}$ is as strict as $Q^{\textrm{strict}}$ and focuses on hyperedges with nodes split into a unique cluster by the partition $\mathcal{C}$. As already stressed, the major difference between $Q^{\textrm{strict}}$ and $Q^{\textrm{aon}}$ lies in that the latter, while summing similar differences as the former, weights differently each terms in those differences (with weights adaptive to the data, as they are estimated from these).
@@ 2.3 Modularity maximization methods
Note that we did not include in our experiments a comparison with methods based on clique reductions. Indeed, Kumar et al. (2020) already did so and concluded that “hypergraph based methods perform consistently better than their clique based equivalents” (end of page 16 in that reference).
@@
Kumar et al. (2020) propose to maximize their modularity $Q^{\textrm{w-clique}}$ relying on the popular and fast Louvain algorithm for graphs (Blondel et al., 2008). More precisely, they do not simply apply Louvain algorithm on the graph $G^{\textrm{w-clique}}$ but rather propose an Iteratively Reweighted Modularity Maximization (IRMM) algorithm where they iteratively apply Louvain on a weighted clique reduction graph, and compute new hyperedge weights (see Algorithm 1 in Kumar et al., 2020). The hyperedge re-weighting step puts a larger weight on hyperedges which are cut into more unbalanced partition vectors by the current partition $\mathcal{C}$. For example, a size-$s$ hyperedge cut into the partition vector $\mathbf{p}=(s-1,1)$ (meaning a unique node falls in a cluster different from the majority one) is much more unbalanced than another one cut into the partition vector $\mathbf{p}=(s/2,s/2)$ (namely half of the nodes belong to a first cluster, the other half belonging to a second cluster) and thus gets a larger weight (see Figure 1 in Kumar et al., 2020). By getting a larger weight, it is more likely that the unique node in this hyperedge will join the majority cluster at Louvain’s next step. The function hmod.kumar implements the IRMM algorithm.
@@
Kamiński et al. (2019a) propose a Clauset-Newman-Moore like (CNM-like) algorithm to maximize $Q^{\textrm{strict}}$ (see Clauset et al., 2004, for the original CNM algorithm). Starting with partition $\mathcal{C}^{\text{own}}$ where each node is in its own part, this algorithm iterates over the set of hyperedges that are split into more than 2 clusters by the current partition, trying to merge all the parts it touches and looking for a modularity improvement. More precisely, the algorithm comes in two versions. In the first one, a loop over all hyperedges is taken, so that at each step all hyperedges are searched and evaluated for merging. In the second one, a stochastic approach is taken which evaluates at each step just one randomly chosen hyperedge (see Algorithm 1 in Kamiński et al., 2019a, for more details).  The stochastic version is computationally less expensive, especially for larger hypergraphs; however it requires to set a maximal number of iterations. In what follows, we choose that second version and set the number of iterations to twice the total number of hyperedges. The implementation is available from Kamiński et al. (2019b), in a mix of Python and Julia files. More precisely, a script strictModularity.py contains a “quick” Python implementation that should work on small datasets only, while a Julia function find_comms is more generally provided to perform the CNM-like algorithm. We rely on the latter in our experiments.
@@ 2.4 Synthetic models for binary and modular hypergraphs
To compare the different modularity-based approaches for clustering hypergraphs nodes, it is mandatory to rely on simulations of modular hypergraphs where ground truth clusters are known. As mentioned earlier, there is no single standard method for generating modular hypergraphs, and, to our knowledge, there are two main approaches. The first approach is based on hypergraph stochastic block models, with several variants proposed in the literature. The second approach involves a generalization of the LFR model for graphs (Lancichinetti et al., 2008). We chose to consider two variants of the first approach and the only one that we are aware of in the second approach. A summary of these models is given in Table 2. We highlight the similarities and differences between those different generating models and the characteristics of the hypergraphs generated by those approaches.
In all those models, we fix a number of nodes $n$, either fix or randomly generate a true number of clusters $K$ (that might depend on $n$) as well as a true partition of the nodes $\mathcal{C}^{\text{true}}=(C_{1}^{\text{true}},\dots,C_{K}^{\text{true}})$ and a maximal size $S$ of hyperedges.
@@ 3.1 Scenarios
We start with “base case” scenarios, called scenarios A, that we defined in the 3 different generating models (HSBM, DCHSBM and h-ABCD). Then we explore other scenarios (called B to F and Z), simulated under the most convenient model to do so, and in which we modify only one characteristic at a time wrt the base case. Each scenario is composed of sub-cases with different samples sizes, comprising in general cases 1 to 6 corresponding to $n\in\{50,100,150,200,500,1000\}$. Moreover, in each scenario explored, we randomly generated 25 hypergraphs. Table 3 gives a summary of the scenarios considered and the empirical characteristics of the 25 hypergraphs generated for each of them. In this table, each simulation is summarized through its differing characteristic wrt the base case (namely, scenarios A). For example, ScenB-DCHSBM is a simulation of hypergraphs less sparse than the base case.
@@
We started from a first series of scenarios, called scenarios A, that play the role of a reasonable sparse case for the methods to work. To explore the robustness of our conclusions, these scenarios are presented under the 3 different generating models (HSBM, DCHSBM and h-ABCD) relying on similar settings for sample size $n$ and number of hyperedges $|\mathcal{E}|$. We set the numbers of hyperedges such that they grow linearly with the number of nodes $n$ (sparse setting). We generated $K=3$ clusters with equal size or probability (depending on the generating model) and the maximum hyperedge size $S=3$. This latter choice ensures both reasonable computing times and simplicity of model parametrization. The ratio $|\mathcal{E}_{2}|/|\mathcal{E}|=|\mathcal{E}_{2}|/(|\mathcal{E}_{2}|+|\mathcal{E}_{3}|)$ is set to 0.7 (on average) to reflect the fact that we expect larger sizes hyperedges to be less frequent than smaller-sizes ones. The within-cluster over between-cluster hyperedge ratio is constant wrt size $s\in\{2,3\}$ and set to $\rho_{s}=1.7$ (either exactly or on average), in order to obtain modular hypergraphs.